OpenStudy (anonymous):

Getting nowhere... Calculate f'(pi / 3) when f(x) = 2 sin x

hartnn (hartnn):

could you differentiate 2sin x ? hint derivative of sin x = cos x

OpenStudy (anonymous):

hmm, I thought you had to convert f'(pi / 3) to a primitive function?

OpenStudy (anonymous):

If I derivate "the right side" I get:

hartnn (hartnn):

i don't know what a primitive function is... f'(x) means derivative of f(x) f'(pi/3) means derivative of f(x) at x = pi/3

OpenStudy (anonymous):

If I derivate "the right side" I get: pi / 3 = 2 cos x ...is that solvable?

OpenStudy (anonymous):

oh and I have to do it without a calculator, so it *should* be something simple :P

OpenStudy (anonymous):

well, nothing, I have no idea how to do that calculation without a calculator, which is why I think it's wrong. I think pi / 3 should be converted to a primitive function and *that* value set to 2 sin x

hartnn (hartnn):

no, f'(pi/3) means derivative of f(x) at x = pi/3 so, put x=pi/3 in f'(x) f'(pi/3) = 2 cos (pi/3)=...?

hartnn (hartnn):

cos (pi/3) can be found without calculator, as pi/3 is standard angle

OpenStudy (anonymous):

hmm, you're right!

hartnn (hartnn):

i know! :')